# Multi-mode Coherent Detection Ghost Imaging Lidar and Vibration-Mode Imaging

### v1

@@ 2.3 Analysis
The previous section presents a comprehensive analytical framework for the bucket-detector-based multi-mode CD-GI lidar system. In the following, we proceed with a detailed analysis.
The imaging receiving module employed in this work constitutes a imaging system, as described in Eq. (4). Under the paraxial approximation and a fixed receiving optical configuration, the maximum angular FoV satisfies $\theta_{\mathrm{FoV}}\approx D_{\mathrm{det}}/f$, where $D_{\mathrm{det}}$ denotes the effective detector width and $f$ is the focal length of the receiving lens. Accordingly, the FoV on the target plane scales linearly with the detector size and can be expressed as $\mathrm{FoV}\approx\theta_{\mathrm{FoV}}\cdot z_{3}\approx D_{\mathrm{det}}\cdot M$, where $M=z_{3}/z_{4}$ denotes the magnification of the imaging receiving module. Therefore, under the assumption of an unbounded image-side FoV, for a fixed receiving optical system, enlarging the detector directly expands the accessible imaging area on the target plane.
2.3.1 Analysis for the IF signal energy of reflected light fields
Under the paraxial approximation, the IF signal in Eq. (6) can be reformulated in the wave-vector domain by expressing both the reflected multi-mode light fields and LO fields using angular spectrum representation[54, 55] and integrating over the finite detector area. The detailed derivation is provided in Supplement 1, Sec. S3.A. The resulting expression can be written as:
@@ 2.3.1 Analysis for the IF signal energy of reflected light fields
where $\kappa_{c}=\frac{D_{f}}{\lambda z_{4}}$ is the receiving system cutoff spatial frequency, $\boldsymbol{\kappa}_{\mathrm{lo}}$ denotes the transverse wave vector (angular spectrum component) of the single-mode LO and $D_{\mathrm{det}}$ is the diameter of bucket detector. This integral expression with $\boldsymbol{\kappa}$ explicitly shows that the detector area imposes a spatial-frequency weighting on the IF signal, thereby determining whether the detection process operates in the single-mode or the multi-mode regime.
@@
    Single-mode light field: When the detector diameter $D_{\mathrm{det}}$ is smaller than the effective Airy disk corresponding to the receiving aperture, the detector can be approximated as a point detector, yielding $D_{\mathrm{det}}\leq 2\lambda z_{4}/D_{f}$. In this case, the detected field can be regarded as single-mode. As the detector size continues to increase, the IF signal energy exhibits a tendency toward stable oscillations, as dictated by the integral properties of the Bessel function and consistent with the Siegman antenna theorem[56].
@@
    Multi-mode light field: When the detector diameter exceeds the Airy disk $D_{\mathrm{det}}>2\lambda z_{4}/D_{f}$, the detector behaves as a bucket detector that integrates multiple spatial modes over its active area. In this regime, the detected reflected light fields must be treated as multi-mode. With further increase in detector size, the integration includes multiple single-mode fields, and the IF signal energy statistically converges to a stable oscillatory regime as described by Siegman antenna theorem[56].
In summary, the coherent detection performance is governed by the spatial-mode matching between the reflected light fields and LO fields, together with the detector’s effective area, as described by Siegman antenna theorem. A small detector favors single-mode detection, whereas increasing the detector size enables the collection of multiple spatial modes. In the presence of reflected multi-mode light fields, the IF signal energy rises to saturation and then oscillates as described in Siegman antenna theorem.
2.3.2 Analysis for the energy of field correlation image
The field correlation can be expressed in the wave-vector domain as:
$G[\boldsymbol{\rho}_{r}^{(m,n)}]$ $\propto\sum_{m,n}\Bigg\{\tilde{t}_{o}\!\left(\alpha\,\boldsymbol{\kappa}_{m,n}\right)*\left[\mathrm{sinc}\!\left(\beta\,\boldsymbol{\kappa}_{m,n}\right)e^{j2\pi\gamma\,\boldsymbol{\rho}_{r}^{(m,n)}\!\cdot\!\boldsymbol{\kappa}_{m,n}}\right]\Bigg\}$
$\quad\times\Delta\kappa^{2}\,\prod\!\left(\frac{|\boldsymbol{\kappa}_{m,n}|}{\kappa_{c}}\right)\pi\left(\frac{D_{\mathrm{det}}}{2}\right)^{2}\frac{2J_{1}\!\left(\tfrac{D_{\mathrm{det}}}{2}|\boldsymbol{\kappa}_{m,n}-\boldsymbol{\kappa}_{\mathrm{lo}}|\right)}{\tfrac{D_{\mathrm{det}}}{2}|\boldsymbol{\kappa}_{m,n}-\boldsymbol{\kappa}_{\mathrm{lo}}|},$ (12)
@@ 2.3.2 Analysis for the energy of field correlation image
where $\tilde{t}(\boldsymbol{\rho}_{o})=\exp\left[\frac{j\pi}{\lambda z_{2}}|\boldsymbol{\rho}_{o}|^{2}+\frac{j\pi z_{1}}{\lambda z_{2}^{2}}|\boldsymbol{\rho}_{o}|^{2}+\frac{j\pi}{\lambda z_{3}}|\boldsymbol{\rho}_{o}|^{2}\right]\tilde{T}(\boldsymbol{\rho}_{o})$ denotes the effective target response, defined as the product of the target’s complex reflectivity and the quadratic phase factor introduced by free-space propagation, $\boldsymbol{\rho}_{r}^{(m,n)}$ indicates the discretized reference arm coordinates at the index $(m,n)$ and $\boldsymbol{\kappa}_{m,n}$ denotes the discretized wave vector at the index $(m,n)$. The detailed derivation is provided in Supplement 1, Sec. S3.B.
Assuming the minimum feature size of the pattern $E_{s}^{(i)}(\boldsymbol{\rho}_{r})$ is $d_{\text{pixel}}$ (e.g., the smallest unit in Hadamard or binary coding), then under far-field conditions, each minimum pattern unit corresponds to an angular spectrum component satisfying: $\boldsymbol{\kappa}_{m,n}\sim z_{4}/(z_{3}d_{\text{pixel}}^{(m,n)})$.
As can be seen from Eqs.(11) and (12), although the IF signal energy of the bucket-detector-based multi-mode CD-GI lidar is constrained by Siegman antenna theorem, by exploiting field correlation, it is able to decouple the reflected multi-mode light fields collected by the bucket detector into effective single-mode components determined by the pixel size, thereby circumventing the constraint imposed by Siegman antenna theorem. In addition, imaging can be achieved using a single bucket detector with a spatiotemporal bandwidth product that is reduced compared with that required by a coherent focal-plane array imaging lidar under the same conditions.
@@ 3 Experiment
The experimental system for the bucket-detector-based multi-mode CD-GI lidar is illustrated in Fig. 2. A narrow-linewidth continuous-wave seed source with a central wavelength of 1550 nm (Shanghai Precilasers Technologies Co. Ltd., FL-SF-1550-S) is transmitted via a single-mode optical fiber to a IQ modulator (MXIQER-LN-30, iXblue). The chirped RF signal driving the modulator is generated by a direct digital synthesizer (Analog Devices Inc., AD9914). The modulation signal has a bandwidth of 800 MHz and a modulation period of 1 ms.
@@
After modulation, the temporal modulated light is divided by a fiber optic splitter (Thorlab, TN1550R1A1, 99:1) into two paths. The weak branch (1%) is collimated using a fiber collimator (Thorlab, F260APC-1550) and used as the LO beam. The stronger branch (99%) is sent through an AOM (Qingjin-OE, G-1550-40-L-B-T-AA-A-Y-L), with a frequency offset of 40MHz, then amplified by an erbium-doped fiber amplifier (Shanghai Precilasers Technologies Co. Ltd.) and used as the object arm. The object arm is transmitted onto DMD (Texas Instruments, DLP-650-L-NIR) for spatial modulation through a fiber collimator(Thorlab, F260APC-1550), and then transmitted onto the target through a simple single-lens imaging system with a focal length of 800 mm and an aperture of 150 mm. Meanwhile, the light field reflected from the target is optically mixed with the LO beam in free space through the spatial optical mixer implemented in the imaging receiving system, which has a focal length of 250 mm and an aperture of 150 mm. The resulting IF signal is then obtained using a free-space balanced detector (Thorlabs, PDB230C).
3.1 Experiments of static target imaging in the bucket-detector-based multi-mode CD-GI lidar
[…]
To investigate the imaging performance of the bucket-detector-based multi-mode CD-GI lidar under different surface scattering characteristics, experiments were conducted using both rough and smooth targets. The overall size of the targets was 55 mm. For the rough target, the object consisted of the four letters “SIOM”, where the letter were fabricated using a high-reflectivity film, while the background region was composed of low-reflectivity black cardboard. For the smooth target, the letter were made of black cardboard, and the background was a planar reflective mirror, forming a smooth reflective surface.
Figure 3: The reconstructed image of amplitude and phase reconstructions for rough and smooth targets. (1.a) Amplitude and (1.b) phase images of the rough target, whose ground truth object is shown in (1.c). (2.a) Amplitude and (2.b) phase images of the smooth target, with the corresponding object shown in (2.c). The smooth target exhibits a clear deterministic phase distribution, whereas the rough target produces a spatially random phase. (2.d) One-dimensional phase profile extracted from the highlighted region in (2.b).
@@ 3.1 Experiments of static target imaging in the bucket-detector-based multi-mode CD-GI lidar
In contrast, for the smooth target, both the amplitude image in Fig. 3(2.a) and the phase image in Fig. 3(2.b) show well-defined spatial structures. In particular, the phase image shows concentric fringe figures characteristic from a smooth surface. To further validate the phase reconstruction, Fig. 3(2.b) and Fig. 3(2.d) show that the experimental results agree well with $\exp\left[\frac{j\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|-\frac{z_{2}}{z_{1}}\boldsymbol{\rho}_{r}|^{2}\right]$ in the theoretical formula, which is caused by both the simple single-lens imaging system and the imaging receiving system.
@@
3.2 Experiments of IF signal energy and reconstructed image energy in the bucket-detector-based multi-mode CD-GI lidar system

### v2

@@ 2.3 Analysis
For the imaging receiving module, the maximum angular FoV is governed by $\theta_{\mathrm{FoV}}\approx D_{\mathrm{det}}/f$, where $D_{\mathrm{det}}$ designates the effective detector width and $f$ represents the focal length of the receiving lens. Consequently, the linear FoV on the target plane scales proportionally with the detector size and can be formulated as $\mathrm{FoV}\approx\theta_{\mathrm{FoV}}\cdot z_{3}\approx D_{\mathrm{det}}\cdot M$, where $M=z_{3}/z_{4}$ characterizes the magnification of the receiving module. It follows that increasing the detector size directly expands the accessible imaging region on the target plane.
2.3.1 Analysis for the IF signal energy
The IF signal can be reformulated in the wave-vector domain by expressing both the LO fields and the reflected multi-mode light fields as described in Eqs.(2) and (4) using angular spectrum representation[55, 56]. The detailed derivation is provided in Supplement 1, Sec. S3.A. The resulting expression can be indicated as:
@@ 2.3.1 Analysis for the IF signal energy
$\prod(\frac{|\boldsymbol{\kappa}|}{\kappa_{c}})=\begin{cases}1,&|\boldsymbol{\kappa}|\leq\kappa_{c},\\[6.0pt] 0,&\text{otherwise}.\end{cases}\qquad\kappa_{c}=\frac{D_{f}}{\lambda z_{4}}$ (12.1)
where $\kappa_{c}=\frac{D_{f}}{\lambda z_{4}}$ represents the cutoff spatial frequency of the imaging receiving module; $\boldsymbol{\kappa}_{\mathrm{lo}}$ denotes the transverse wave vector (angular spectrum component) of the single-mode LO; and $D_{\mathrm{det}}$ is the diameter of the bucket detector. This integral expression in terms of $\boldsymbol{\kappa}$ explicitly shows that the detector area exerts a spatial-frequency weighting on the IF signal, thereby dictating whether the detection process operates in the single-mode or the multi-mode regime.
@@
    Single-mode light field: When $D_{\mathrm{det}}\leq 2\lambda z_{4}/D_{f}$, Eq.(12) can be simplified as:
    $\widetilde{i}^{{(i)}}(f=f_{if})\propto\int_{0}^{\kappa_{c}}\kappa\,\prod\left(\frac{\kappa}{\kappa_{c}}\right)\pi\left(\frac{D_{\mathrm{det}}}{2}\right)^{2}d\kappa.$ (13)
    In this case, the detector diameter is smaller than the diffraction-limited Airy disk, enabling the detector to be approximated as a point detector capturing only a single-mode field. As the detector size increases, the energy of the IF signal scales accordingly as described in Eq. (13), which is consistent with the Siegman antenna theorem [57].
@@
    Multi-mode light field: When $D_{\mathrm{det}}>2\lambda z_{4}/D_{f}$, Eq.(12) is simplified as:
    $\widetilde{i}^{{(i)}}(f=f_{if})\propto\int_{0}^{\frac{\pi}{2}}\int_{0}^{\kappa_{c}}\kappa\,\prod\left(\frac{\kappa}{\kappa_{c}}\right)\frac{2J_{1}\!\big(\tfrac{D_{\mathrm{det}}}{2}\kappa\theta\big)}{\tfrac{D_{\mathrm{det}}}{2}\kappa\theta}d\kappa d\theta.$ (14)
    where $\theta$ denotes the angle between the wave-vectors of the LO field and the reflected multi-mode light fields. As the detector diameter exceeds the Airy disk, it functions as a bucket detector that integrates multiple spatial modes. Consequently, the IF signal energy statistically saturates into a stable oscillatory regime, further validating the predictions of the Siegman antenna theorem [57].
In summary, the coherent detection performance is governed by the spatial-mode matching between the reflected light fields and the LO fields, contingent upon the detector’s effective size, as dictated by the Siegman antenna theorem. While a constrained detector size is conducive to single-mode detection, expanding the detector size facilitates the collection of multiple spatial modes. In the presence of spatially complex reflected multi-mode fields, the IF signal energy initially increases toward saturation and subsequently enters an oscillatory regime, a behavior consistent with the theoretical predictions of the Siegman antenna theorem.
2.3.2 Analysis for the energy of reconstructed image by field correlation
The field correlation can be formulated in the wave-vector domain as:
$G[\boldsymbol{\rho}_{r}^{(m,n)}]$ $\propto\sum_{m^{{}^{\prime}},n^{{}^{\prime}}}\Bigg\{\tilde{t}_{o}\!\left(\alpha\,\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}\right)*\left[\mathrm{sinc}\!\left(\beta\,\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}\right)e^{j2\pi\gamma\,\boldsymbol{\rho}_{r}^{(m,n)}\!\cdot\!\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}}\right]\Bigg\}$
$\times\Delta\kappa^{2}\,\prod\!\left(\frac{|\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}|}{\kappa_{c}}\right)\pi\left(\frac{D_{\mathrm{det}}}{2}\right)^{2}\frac{2J_{1}\!\left(\tfrac{D_{\mathrm{det}}}{2}|\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}-\boldsymbol{\kappa}_{\mathrm{lo}}|\right)}{\tfrac{D_{\mathrm{det}}}{2}|\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}-\boldsymbol{\kappa}_{\mathrm{lo}}|},$ (15)
@@ 2.3.2 Analysis for the energy of reconstructed image by field correlation
where $\tilde{t}(\boldsymbol{\rho}_{o})=\exp\left[\frac{j\pi}{\lambda z_{2}}|\boldsymbol{\rho}_{o}|^{2}+\frac{j\pi z_{1}}{\lambda z_{2}^{2}}|\boldsymbol{\rho}_{o}|^{2}+\frac{j\pi}{\lambda z_{3}}|\boldsymbol{\rho}_{o}|^{2}\right]\tilde{T}(\boldsymbol{\rho}_{o})$ represents the effective target response, defined as the product of the target’s complex reflectivity and the quadratic phase factors arising from free-space propagation. Furthermore, $\boldsymbol{\rho}_{r}^{(m,n)}$ designates the discretized reference arm coordinates at index $(m,n)$, while $\boldsymbol{\kappa}_{m^{{}^{\prime}},n^{{}^{\prime}}}$ denotes the discretized wave vector at index $(m^{{}^{\prime}},n^{{}^{\prime}})$. A detailed derivation is provided in Supplement 1, Sec. S3.B. Upon retrieving the complex reflectivity distribution, the total image energy is quantified by summing the intensities across all pixels: $\sum_{m,n}|G[\boldsymbol{\rho}_{r}^{(m,n)}]|^{2}$.
As evidenced by Eqs. (12) and (15), the energy of the IF signal is determined by the modulus squared of the integrated multi-mode light field: $|\widetilde{i}^{{(i)}}(f=f_{if})|^{2}$. In contrast, field correlation allows for the decomposition of the IF signal into individual pixels; the image energy is then calculated by taking the modulus squared of each resolved pixel and summing the results: $\sum_{m,n}|G[\boldsymbol{\rho}_{r}^{(m,n)}]|^{2}$. Crucially, while the IF signal energy of the bucket-detector-based CD-GI lidar remains constrained by the Siegman antenna theorem, the application of field correlation decouples the multi-mode fields.
This process effectively partitions the collective IF signal into single-mode components defined by the pixel resolution, thereby circumventing the limitations in the Siegman theorem. Moreover, imaging can be achieved using a single bucket detector with a low spatiotemporal bandwidth product.
@@ 3 Experiment
The experimental system for the bucket-detector-based multi-mode CD-GI lidar is illustrated in Fig. 2. A narrow-linewidth continuous-wave seed source with a central wavelength of $1550$ nm (Shanghai Precilasers Technologies Co. Ltd., FL-SF-1550-S) is coupled into a single-mode optical fiber and fed to an IQ modulator (MXIQER-LN-30, iXblue). The chirped RF signal driving the modulator is generated by a direct digital synthesizer (AD9914, Analog Devices Inc.), featuring a modulation bandwidth of $800$ MHz and a period of $1$ ms.
@@
Following modulation, the temporally modulated light is partitioned by a fiber optic splitter (Thorlab, TN1550R1A1, 99:1) into two paths. The $1\%$ branch is collimated (Thorlab, F260APC-1550) to serve as the LO beam. The $99\%$ branch passes through an AOM (Qingjin-OE, G-1550-40-L-B-T-AA-A-Y-L) to introduce a frequency offset of $40$ MHz. The beam is subsequently amplified by an erbium-doped fiber amplifier (Shanghai Precilasers Technologies Co. Ltd.) and spatially encoded via a DMD (Texas Instruments, DLP-650-L-NIR). The modulated beam is transmitted onto the target through a simple single-lens imaging module with a focal length of 800 mm and an aperture of 150 mm. The light field reflected from the target is optically mixed with the LO beam in free space via the spatial optical mixer implemented in the imaging receiving module, which has a focal length of 250 mm and an aperture of 50.8 mm. The resulting IF signal is then obtained using a free-space balanced detector (Thorlabs, PDB230C), which the maximum photosensitive area diameter is 0.3 mm.
3.1 Experiments of static target imaging with the bucket-detector-based multi-mode CD-GI lidar
In the static target imaging experiments, the imaging transmission module in Fig. 2 yields a minimum resolvable feature size of approximately $1.22\lambda f/D\approx 10~\mu\mathrm{m}$. The DMD has a resolution of $1280\times 800$ with $10.8~\mu\mathrm{m}$ micromirrors and operates at a refresh rate of 4000 Hz. To match the resolution of imaging transmission module, the minimum unit of Hadamard pattern is set to a 1-pixel binning, resulting in an encoded block size of $10.8~\mu\mathrm{m}$. The effective coding region is $64\times 64$ modulation array after binning. Under the experimental conditions at 65 m, the emission angle is 0.85 mrad, corresponding to the FoV of 55.5 mm.
To investigate the imaging performance of the bucket-detector-based multi-mode CD-GI lidar under different surface roughnesses of targets, experiments were conducted using both rough and smooth targets. The overall size of the targets was 55 mm. For rough target, the object consisted of the four letters “SIOM”, where the letter were fabricated using a high-reflectivity film, while the background region was composed of low-reflectivity black cardboard. For smooth target, the letter were made of black cardboard, and the background was a planar reflective mirror, forming a smooth reflective surface.
Figure 3: The reconstructed image of amplitude and phase reconstructions for rough and smooth targets. (1.a) Amplitude and (1.b) phase images of the rough target, whose ground truth object is shown in (1.c). (2.a) Amplitude and (2.b) phase images of the smooth target, with the corresponding object shown in (2.c). The smooth target exhibits a clear deterministic phase distribution, whereas the rough target produces a spatially random phase. (2.d) Comparison of the theoretical curve and experimental results of the one-dimensional phase distribution.
@@ 3.1 Experiments of static target imaging with the bucket-detector-based multi-mode CD-GI lidar
In contrast, for the smooth target, both the amplitude image in Fig. 3(2.a) and the phase image in Fig. 3(2.b) show well-defined spatial structures. In particular, the phase image shows concentric fringe figures characteristic. To further validate the phase reconstruction, Fig. 3(2.b) and Fig. 3(2.d) show that the experimental results agree well with the term $\exp\left[\frac{j\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|-\frac{z_{2}}{z_{1}}\boldsymbol{\rho}_{r}|^{2}\right]$ in Eq.(12), which is caused by both the simple single-lens imaging transmission module and the imaging receiving module.
@@
3.2 Experiments of IF signal energy and reconstructed image energy with the bucket-detector-based multi-mode CD-GI lidar system
As shown in Eq. (12), to investigate the relationship among the imaging receiver aperture, detector area, and the IF signal energy, a series of black pinhole masks with diameters of 0.1 mm, 0.2 mm, and 0.3 mm (Edmund Optics) were sequentially placed on the detector plane to control the effective detection area. These pinhole diameters corresponded to FoV diameters of 25.9 mm, 51.8 mm, and 77.7 mm on the target plane at 65 m. The minimum unit of Hadamard pattern is set to a 4-pixels binning. The effective coding region is $64\times 64$ modulation array after binning. To regulate the receiving aperture, a variable diaphragm is positioned before the lens within the imaging receiving module. The target was a circular reflectivity plate with a diameter of 100 mm, ensuring uniform filling of the maximum effective detection area of the detector. The IF signal energy was obtained from the same number of repeated measurements of the first fully illuminated pattern followed by statistical averaging, while the reconstructed image energy was computed by converting the reconstructed intensity into relative energy values. The pixel count of the target area is obtained by normalizing the reconstructed image to its maximum intensity and applying a 2% intensity threshold (0.02) to classify and isolate the target pixels.
We confirmed a clear matching relationship between the receiver aperture and detector size that maximizes the IF signal energy, as described by Eqs. (13) and  (14). The experimental results are shown in Fig. 4 (a). Under single-mode LO conditions, the receiving apertures for detectors with diameters of 0.3 mm, 0.2 mm, and 0.1 mm were measured to be 3.00 mm, 4.50 mm, and 9.50 mm, closely matching the theoretical predictions described by Eqs. (13) and  (14) of 3.15 mm, 4.73 mm, and 9.46 mm. This result is also in agreement with the effective mode field matching conditions predicted by the Siegman antenna theorem[57].
Figure 4: Normalized IF signal energy and the relative energy ratio of the reconstructed images versus the receiving aperture variation range under different bucket detector diameters (simulation and experiment). (a) Normalized IF signal energy as a function of the receiving aperture variation range for various bucket detector diameters: blue represents a detector diameter of 0.1 mm; green represents a detector diameter of 0.2 mm; red represents a detector diameter of 0.3 mm; the circle represents the experimental curve; the asterisk represents the simulation curve. The horizontal axis represents the receiving aperture, and the vertical axis represents the normalized energy. (b) Relative energy ratio of the reconstructed images as a function of the bucket detector diameter: the asterisk denotes the simulation curve, while the blue solid line represents the experimental curve. The horizontal axis represents the bucket detector diameter, and the vertical axis represents the ratio of relative energy.
Furthermore, the statistical characteristics of the IF signal energy in the bucket-detector-based multi-mode CD-GI lidar system and its relationship with the total energy of the reconstructed image by field correlation have been systematically evaluated. The experimental results in Fig. 4 (b) show that the ratio of the total energy of all reconstructed image pixels to the IF signal energy varies with detector size. Specifically, the measured ratios were 680.9 for $D_{\mathrm{det}}=0.3$ mm, 271.6 for $D_{\mathrm{det}}=0.2$ mm and 98.5 for $D_{\mathrm{det}}=0.1$ mm. These ratios closely match the theoretical expectations based on the effective number of pixels shown in Eq. (15) (697, 300 and 78, respectively, as detailed in Supplement 1, Sec. S4.A), directly validating the ability to separate the sum of the reflected multi-mode light fields to pixels using field correlation.
Experiments on the IF signal energy and image energy from field-correlation reconstruction jointly demonstrate that, although the IF signal energy is constrained by Siegman antenna theorem, the exploitation of field correlation enables the reflected multi-mode light fields collected by the bucket detector to be coherently decoupled into effective single-mode components determined by the pixel size, thereby effectively circumventing the constraint imposed by Siegman antenna theorem.
3.3 Experiments of vibration modes imaging
The bucket-detector-based multi-mode CD-GI lidar is further applied in experiments to assess its capability in vibration modes imaging of the target.
3.3.1 Experiments of one vibrating target:
A loudspeaker (8 $\Omega$, 5 W) was employed as the vibration source. Because the diaphragm material of the speaker has very low reflectivity at 1550 nm, a high-reflectivity (HR) film was affixed to its surface to enhance the optical reflectivity. The target reflectivity distribution was set as illustrated in Fig. 5(a), featuring a rectangular geometry with a length of $60$ mm and a width of $25$ mm. This specific target scale was chosen to ensure it remains well within the detection FoV at the $65$ m. The loudspeaker was driven by a sinusoidal voltage signal generated from a computer sound card, with vibration frequencies set at 1000 Hz and 2000 Hz. To capture as much as possible the entire vibration cycle and record the complete vibration state of the target, the repetition frequency of the DMD is set at $1000$ Hz. The effective coding region is $64\times 64$ modulation array after 2-pixels binning (Hadamard patterns). The detection FoV is measured at $77$ mm. The sampling rate is established at $125$ MHz to satisfy the Nyquist sampling theorem for the $40$ MHz IF signal.
For each pattern acquisition, $110,016$ data points are collected, corresponding to a total duration of $880~\mu\text{s}$. The window length for the STFT is defined as $4096$ points (approximately $33~\mu\text{s}$). Within this localized time interval, the target motion is approximated as uniform velocity (see Supplement 1, Sec. S2.B for further details). An overlap length of $2048$ points (approximately $17~\mu\text{s}$) is applied between adjacent windows to ensure temporal continuity.
Table 1: Key Experimental and Signal Processing Parameters
Parameter Value Unit
DMD Repetition Frequency 1000 Hz
Sampling Rate 125 MHz
IF Signal Frequency 40 MHz
Points per Each Pattern 110,016 -